OpenStudy (anonymous):

Find the sum of the first 11 terms in the sequence 4,6,9...

OpenStudy (anonymous):

Do you know how to do it and just want the answer? or do you need to know how to do it....? lol

OpenStudy (anonymous):

13

OpenStudy (anonymous):

yea that haha show off >:)

OpenStudy (anonymous):

need to know how to do it

OpenStudy (anonymous):

yay!! (:

OpenStudy (anonymous):

haha why dont i let mr ivan whatever tell u!! :P

OpenStudy (anonymous):

see...

OpenStudy (anonymous):

it's pretty easy!

OpenStudy (anonymous):

4+2=6 6+3=9 9+4=13

OpenStudy (anonymous):

o i see i van! thanks

OpenStudy (anonymous):

haha i dont even have to do any work just let this guy do it!! :O

OpenStudy (radar):

396

OpenStudy (anonymous):

howd u get that radar

OpenStudy (radar):

Your original problem wanted the sum of the first 11 terms of the sequence. Ivanplayer showed the sequence, just add them up. I actually added up 11 terms starting with 6, instead of 4. You need to subtract 77 from that. The first 11 terms are: 4,6,9,13,18,24,31,39,48,58,69, for a total of 319